1. Overhead Variances, Four-Variance Analysis

    Oerstman, Inc., uses a standard costing system and develops its overhead rates from the current annual budget. The budget is based on an expected annual output of 120,000 units requiring 480,000 direct labor hours. (Practical capacity is 500,000 hours.) Annual budgeted overhead costs total $796,800, of which $571,200 is fixed overhead. A total of 119,400 units using 478,000 direct labor hours were produced during the year. Actual variable overhead costs for the year were $260,400, and actual fixed overhead costs were $556,250.

    Required:

    1. Compute the fixed overhead spending and volume variances.

    Fixed Overhead Spending Variance $
    Fixed Overhead Volume Variance $

    2. Compute the variable overhead spending and efficiency variances. Do not round intermediate calculations

    Variable Overhead Spending Variance $
    Variable Overhead Efficiency Variance $
Answer #1

1) Fixed Overhead predetermined rate = Budgeted Fixed Overhead/Budgeted units

= $571,200/120,000 units = $4.76 per unit

Fixed Overhead Spending Variance = Budgeted Fixed Overhead - Actual Fixed Overhead

= $571,200 - $556,250 = $14,950 Favorable

Fixed Overhead Volume Variance = Applied Fixed Overhead - Budgeted Fixed Overhead

Applied Fixed Overhead = Actual units produced*Fixed Overhead predetermined rate

= 119,400*$4.76 per unit = $568,344

Fixed Overhead Volume Variance = $568,344 - $571,200 = $2,856 Unfavorable

Therefore, fixed overhead spending variance is $14,950 Favorable and fixed overhead volume variance is $2,856 Unfavorable.

2) Budgeted labor hours per unit = Budgeted Labor hours/Budgeted units

= 480,000 hrs/120,000 units = 4 hours per unit

Std labor hours for actual units = Actual units*budgeted labor hour per unit

= 119,400 units*4 hrs per unit = 477,600 hours

Budgeted variable overhead = $796,800 - $571,200 = $225,600

Budgeted variable overhead rate per hour = $225,600/480,000 hours = $0.47 per hour

Variable Overhead Spending Variance

= (Budgeted variable OH rate*Actual direct labor hours) - Actual variable overheads

= ($0.47*478,000 hrs) - $260,400 = 35,740 Unfavorable

Variable Overhead Efficiency Variance = (Std labor hrs - Actual labor hrs)*Budgeted variable OH rate

= (477,600 - 478,000)*$0.47 = $188 Unfavorable

Therefore, variable overhead spending variance is $35,740 Unfavorable and variable overhead efficiency variance is $188 Unfavorable.